A vehicle moving with a constant acceleration from A to B in a straight line AB, has velocities u and v at A and B. C is the mid point of line AB. If the time taken to travel from A to C is twice the time taken to travel from C to B then the velocity of the vehicle at B is ? 

1) 5u    2) 6u   3) 7u    4) 8u

Kindly please ,explain me this problem im stuck with .! Thank you :)

Grade:12

1 Answers

noogler
8 years ago
velocity at B is 7u
                                 t1                                      t2
                .-----------------------------------------.----------------------------------------.                      t1=2t2
              A  (u)                                    C (vl)                                B(v)
 
from A to C:                                                   from C to B:
 
vl=u+at1                                                                v=vl+at2
v1=u+2at2                                                             vl=v-at2
                               from above eq.  u-v = -3at …........................(1)
x=ut1+1/2at12                                                                              x=vlt2+1/2at22        substitute v1=v-at2 in it 
                              solving aove two eq. u get 2u-v= -5/2at2  …............................(2)
 
 
solving 1 and 2 u get                 u=1/2at2            sub in u-v= -3at2              v=7/2at2................v=7u
 
 
APPROVE IF USEFUL!!!!!!!!!!!!!!!!!!!